Circular No. 8105

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003cx Further to IAUC 8104, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of an apparent supernova (mag 19.1) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Mar. 31.42, 31.46, and 31.48 UT. The new object is located at α = 13^{h}57^{m}06^{s}.46, δ = −17^{o}02′22′′.6 (equinox 2000.0), which is 1′′.0 west and 2′′.6 south of the center of the apparent host galaxy. Additional magnitudes from NEAT observations with the Palomar 1.2-m and Haleakala 1.2-m telescopes: 2002 May 17.29, [21.0; 2003 Feb. 12.61, [20.3; Mar. 23.44, 19.4.

SUPERNOVA 2003cp IN MCG +10-12-78 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003cp (cf. IAUC 8103), obtained by P. Berlind on Apr. 1.18 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to resemble a type-Ia supernova, two to three weeks past maximum, but some of the line strengths appear unusual.

IGR J17464-3213 = XTE J17464-3213 M. P. Rupen, A. J. Mioduszewski, and V. Dhawan, National Radio

Astronomy Observatory (NRAO), report the detection of a probable radio counterpart to the the x-ray transient and blackhole candidate IGR J17464-3213 = XTE J17464-3213 (Revnivtsev et al. 2003, http://atel.caltech.

edu/?read=132; Markwardt and Swank 2003, http://atel.caltech.edu/

?read=133). Observations with the Very Large Array (VLA) on Mar. 30 and Apr. 1 UT show a compact, variable source at α = 17^{h}46^{m}15^{s}.61 ± 0^{s}.01, δ = −32^{o}13′59′′.9 ± 1′′ (equinox 2000.0), roughly 2′ from the original INTEGRAL position and consistent with the position of H1743–322 (Gursky et al. 1978, Ap.J. 223, 973). The flux densities on Mar. 30 (after correcting for primary beam attenuation) were 3.5 ± 0.2 mJy at 4.9 GHz and 4.6 ± 0.2 mJy at 8.5 GHz. By Apr. 1, the source had brightened by ∼ 50 percent, while retaining a flat or somewhat rising spectrum: 1.4 GHz, 7.8 ± 1.5 mJy; 4.9 GHz, 6.0 ± 0.3 mJy; 8.5 GHz, 6.6 ± 0.2 mJy; 14.9 GHz, 6.9 ± 0.6 mJy. The NRAO VLA Sky Survey (NVSS; Condon et al. 1998, A.J. 115, 1693) shows no source at 1.4 GHz in 1995, in an image with an rms noise of 0.45 mJy/beam. Further radio observations are planned, and observations at other wavelengths are urgently requested.

2003 April 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green